# ChunkRAG: Novel LLM-Chunk Filtering Method for RAG Systems

### v3

@@ 5.1 Evaluation Metrics
We used accuracy as the primary evaluation metric, calculated as the percentage of generated responses that exactly match the ground-truth answers. Additionally, we considered semantic accuracy, where responses that are semantically equivalent to the ground truth (as evaluated by an LLM-based semantic similarity checker) are also counted as correct.
Accuracy=Correct  AnswersTotal  $\text{Accuracy}=\frac{\text{Correct \ Answers}}{\text{Total \ Questions}}$ (3)
@@ 5.2 Comparison and Impact
While a 10 percentage point increase may seem incremental, it translates into an exponential improvement in output effectiveness in practical applications. This is particularly evident when considering the error rates and their impact on the overall user experience.
5.3 Multi-Step Processes
In applications requiring multi-hop reasoning or sequential decision-making, errors can compound exponentially. The probability of the system providing correct answers across multiple steps is given by:
$P(\text{correct \ over }\,n\,\text{ steps})=(\text{accuracy})^{n}$ (4)
As the number of steps increases, the gap between the success probabilities widens exponentially. For a 3-step process, our method’s success rate is 66% higher than CRAG’s. This exponential improvement is especially important in complex tasks where each additional step compounds the risk of error, namely relevant to OpenAI’s advanced models such as o1 where the language model utilizes multi-hop reasoning, relying on spending time "thinking" before it answers, making it more efficient in complex reasoning tasks, science and programming.

### v4

@@ 5.1 Evaluation Metrics
We used accuracy as the primary evaluation metric, calculated as the percentage of generated responses that match the ground-truth answers.
@@ 5.2 Comparison and Impact
While a 10 percentage point increase may seem incremental, it translates into an exponential improvement in output effectiveness in practical applications. This is particularly evident when considering the error rates and their impact on the overall user experience.In applications requiring multi-hop reasoning or sequential decision-making, errors can compound exponentially. This exponential improvement is especially important in complex tasks where each additional step compounds the risk of error, namely relevant to OpenAI’s advanced models such as o1 where the language model utilizes multi-hop reasoning, relying on spending time "thinking" before it answers, making it more efficient in complex reasoning tasks, science and programming.